OpenStudy (friendzone):

-2/5 divided (-9/17)

OpenStudy (friendzone):

this one is finding the quotient

Directrix (directrix):

-2/5 divided (-9/17) = -2/5 * 17/(-9) = ?

Directrix (directrix):

Yes, it is a quotient problem. To divide one fraction by another, invert the fraction behind the division symbol and multiply numerators together as well as denominators.

Directrix (directrix):